A clay model of an automobile is 8 inches long and weighs 2 pounds. What would a full size clay model of a similar automobile weigh if it were 12 feet long?
Answer by scott8148(6628)

the weight is proportional to the volume of the model

the volume is proportional to the cube of the length

so the weight is proportional to the cube of the length __ 12 feet is 144 inches

let x=weight __ 2/8^3=x/(144)^3 __ 2(18)^3=x